# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ B.2 Probit Model
Static Model
$\widehat{\beta}$ $0.150$ $0.197$ $0.063$ $0.082$ $0.042$ $0.054$
$\widehat{\beta}_{L}$ $0.055$ $0.124$ $0.013$ $0.049$ $0.007$ $0.032$
Dynamic Model
$\widehat{\beta}$ $0.159$ $0.247$ $0.067$ $0.086$ $0.043$ $0.055$
$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.134$ $0.015$ $0.052$ $0.008$ $0.032$
$\widehat{\beta}_{L}^{(2)}$ $0.062$ $0.136$ $0.016$ $0.053$ $0.008$ $0.032$
$\widehat{\rho}$ $-0.055$ $0.124$ $-0.022$ $0.054$ $-0.017$ $0.036$
$\widehat{\rho}_{L}^{(1)}$ $-0.030$ $0.105$ $-0.009$ $0.047$ $-0.007$ $0.031$
$\widehat{\rho}_{L}^{(2)}$ $-0.022$ $0.106$ $-0.003$ $0.047$ $-0.003$ $0.031$
@@
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table B.7: Empirical Size of LR, LM, and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 1
(a) LR Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $29\%$ $13\%$ $27\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $23\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
(b) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $28\%$ $13\%$ $26\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
(c) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $28\%$ $13\%$ $25\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
@@
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.8: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 1
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $59\%$ $41\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $96\%$ $71\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $83\%$ $63\%$ $40\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $59\%$ $40\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $95\%$ $71\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $83\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $58\%$ $40\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $95\%$ $72\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $95\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $82\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
@@
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Table B.9: Empirical Size of LM and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 2
(a) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $69\%$ $34\%$ $59\%$ $30\%$
$(60,60)$ $74\%$ $18\%$ $62\%$ $17\%$
$(90,90)$ $71\%$ $14\%$ $60\%$ $13\%$
(b) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $37\%$ $29\%$ $25\%$ $26\%$
$(60,60)$ $71\%$ $18\%$ $56\%$ $16\%$
$(90,90)$ $72\%$ $14\%$ $60\%$ $13\%$
@@
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.10: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 2
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $97\%$ $90\%$ $77\%$ $14\%$ $13\%$ $15\%$
$(60,60)$ $100\%$ $100\%$ $96\%$ $34\%$ $70\%$ $83\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $85\%$ $98\%$ $96\%$
Dynamic Model
$(30,30)$ $95\%$ $86\%$ $74\%$ $25\%$ $28\%$ $38\%$
$(60,60)$ $100\%$ $100\%$ $93\%$ $47\%$ $84\%$ $96\%$
$(90,90)$ $100\%$ $100\%$ $100\%$ $90\%$ $99\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $96\%$ $87\%$ $73\%$ $12\%$ $13\%$ $17\%$
$(60,60)$ $100\%$ $100\%$ $96\%$ $37\%$ $69\%$ $76\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $86\%$ $91\%$ $88\%$
Dynamic Model
$(30,30)$ $94\%$ $83\%$ $70\%$ $21\%$ $27\%$ $38\%$
$(60,60)$ $100\%$ $99\%$ $92\%$ $48\%$ $81\%$ $91\%$
$(90,90)$ $100\%$ $99\%$ $99\%$ $90\%$ $94\%$ $94\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $91\%$ $83\%$ $68\%$ $10\%$ $13\%$ $19\%$
$(60,60)$ $99\%$ $99\%$ $95\%$ $39\%$ $76\%$ $92\%$
$(90,90)$ $99\%$ $99\%$ $99\%$ $88\%$ $99\%$ $99\%$
Dynamic Model
$(30,30)$ $88\%$ $75\%$ $60\%$ $21\%$ $27\%$ $39\%$
$(60,60)$ $100\%$ $99\%$ $91\%$ $49\%$ $85\%$ $98\%$
$(90,90)$ $99\%$ $99\%$ $99\%$ $91\%$ $99\%$ $100\%$

### v5

@@ F.2 Heterogeneous Autoregressive Parameters
Static Poisson Model
$\widehat{\beta}$ $0.016$ $0.043$ $0.005$ $0.018$ $0.003$ $0.012$
$\widehat{\beta}_{L}$ 0.002-0.002 $0.039$ 0.002-0.002 $0.018$ 0.001-0.001 $0.011$
$\widehat{\beta}_{J}$ 0.014-0.014 $0.051$ 0.004-0.004 $0.019$ 0.002-0.002 $0.012$
$\widehat{\beta}_{B}$ $0.001$ $0.039$ $0.000$ $0.017$ $0.000$ $0.011$
Dynamic Poisson Model
$\widehat{\beta}$ $0.025$ $0.058$ $0.008$ $0.022$ $0.005$ $0.014$
$\widehat{\beta}_{L}^{(1)}$ 0.005-0.005 $0.051$ 0.004-0.004 $0.021$ 0.002-0.002 $0.014$
$\widehat{\beta}_{L}^{(2)}$ 0.005-0.005 $0.051$ 0.004-0.004 $0.021$ 0.002-0.002 $0.014$
$\widehat{\beta}_{J}$ 0.060-0.060 $0.218$ 0.010-0.010 $0.025$ 0.004-0.004 $0.015$
$\widehat{\beta}_{B}$ 0.003-0.003 $0.054$ $0.000$ $0.021$ $0.000$ $0.014$
$\widehat{\rho}$ 0.204-0.204 $0.250$ 0.067-0.067 $0.075$ 0.041-0.041 $0.045$
$\widehat{\rho}_{L}^{(1)}$ 0.077-0.077 $0.110$ 0.017-0.017 $0.036$ 0.006-0.006 $0.023$
$\widehat{\rho}_{L}^{(2)}$ 0.086-0.086 $0.118$ 0.022-0.022 $0.038$ 0.010-0.010 $0.022$
$\widehat{\rho}_{J}$ $0.588$ $0.821$ $0.060$ $0.083$ $0.014$ $0.027$
$\widehat{\rho}_{B}$ $1.323$ $1.548$ $0.024$ $0.056$ $0.006$ $0.020$
@@
Table F.8: Empirical Size and Power from LR Test for Two-Way Fixed-Effect Probit Estimates
(a) Empirical Size
Size
$(N,T)$ Uncorrected Analytical Bootstrap
$(30,30)$ $53$ $25$ $12$
$(60,60)$ $48$ $12$ 55
$(90,90)$ $46$ 88 44
(b) Empirical Power
Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
$(N,T)$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
$(30,30)$ $99$ $57$ $93$ $100$ $76$ 1919 $56$ $98$
$(60,60)$ $99$ $99$ $99$ $99$ $100$ $89$ $100$ $100$
$(90,90)$ $99$ $99$ $99$ $99$ $99$ $99$ $99$ $99$
@@
For the two binary response models, we find that adding heterogeneities to the autoregressive coefficient may induce some numerical difficulty. To deal with it, we apply certain restrictions to the data-generating process. In Remark 8 below, we first explain the numerical issue, followed by introducing the data-generating process and the restrictions we apply.
Remark 8 (binary response model with heterogeneous autoregressive coefficient)
For binary response models with heterogeneous autogressive coefficient, numerical difficulty may arise if the (lagged) dependent variable contains insufficient variations. Consider a simple binary response model
$Y_{it}=\mathds{1}\{\left(\rho_{0}+\alpha_{1,i}+\gamma_{1,t}\right)Y_{it-1}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}.$
Suppose, for an individual $i$, there are too many time periods $t$ such that $Y_{it-1}=1$. Then, the regressor vector $\left(Y_{i1},\ldots,Y_{iT-1}\right)^{\prime}$ could be nearly collinear to the constant of $1$. This renders the separate identification of $\alpha_{1,i}$ and $\alpha_{2,i}$ numerically difficult. Similar situations also happen if, for some $t$, there are too many individuals with $Y_{it-1}=1$. When there are excessive zeros in $Y_{it-1}$ for some $i$ or $t$, the identification of $\alpha_{1,i}\$or $\gamma_{1,t}$ is also very difficult.
For the simulation, we consider a binary response model specified according to the data-generating process
$Y_{it}=\mathds{1}\{\left(\rho_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0}\right)Y_{it-1}+(\beta_{0}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0})Z_{it}+\delta_{0}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0}+\varepsilon_{it}>0\},$
where, for $k=1,2,3$, $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.04)$ and are demeaned; $\varepsilon_{it}$ are either standard normal (probit) or standard logistic (logit); and $\delta_{0}$ is explained below. All other settings are the same as in Section 4, except that we restrict both $0.1\leq T^{-1}\sum_{t=1}^{T}Y_{it}\leq 0.6$ for all $i$ and $0.1\leq N^{-1}\sum_{i=1}^{N}Y_{it}\leq 0.6$ for all $t$. An additional common parameter $\delta_{0}=-1$ is included in the intercept to make it easier to generate a data set which satisfies the two restrictions [^11] The response $Y_{it}$ tends to contain too many ones without $\delta_{0}=-1$. This is because $\left(\rho_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0}\right)Y_{it-1}$ is positive on average.. Note that $\delta_{0}$ is considered nuisance and is not a part of $\theta_{0}$. These measures seem to alleviate the numerical difficulty considerably, although they are, admittedly, somewhat ad-hoc.
Table F.9 presents the simulation results for both the probit and logit models with heterogeneous autoregressive coefficient. For dynamic models, we see that $\widehat{\rho}$ exhibits larger biases, and our bias correction procedure reduces the bias considerably. Under the current setting, neither the jackknife nor the bootstrap perform better than our bias correction procedure. For static models, the conclusion is the same as in Section 4, except that the biases of all estimators are slightly larger (which is likely caused by having $\delta_{0}=-1$). Table F.10 presents the empirical sizes and powers from the LR tests. We see that the LR test based on the uncorrected likelihood is fatally size-distorted. This is anticipated because the IPP is much more severe with an additional set of heterogeneities. Our bias correction procedure reduces the empirical type-I error risk considerably. Similar to the results in Section 4, the bootstrap reduces the empirical type-I error risk at the cost of reducing the power of the test as well.
Table F.9: Comparisons of Uncorrected and Corrected Probit and Logit Estimates with Heterogeneous Autoregressive Coefficient, Design 1
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Logit Model
$\widehat{\beta}$ $0.196$ $0.255$ $0.081$ $0.097$ $0.050$ $0.059$
$\widehat{\beta}_{L}$ $0.065$ $0.128$ $0.018$ $0.051$ $0.007$ $0.030$
$\widehat{\beta}_{J}$ 1.344-1.344 $1.939$ 0.049-0.049 $0.108$ 0.013-0.013 $0.033$
$\widehat{\beta}_{B}$ 0.143-0.143 $0.179$ 0.014-0.014 $0.046$ 0.006-0.006 $0.029$
Static Probit Model
$\widehat{\beta}$ $0.374$ $0.566$ $0.105$ $0.113$ $0.062$ $0.066$
$\widehat{\beta}_{L}$ $0.125$ $0.164$ $0.038$ $0.053$ $0.017$ $0.027$
$\widehat{\beta}_{J}$ 5.362-5.362 $7.774$ 0.350-0.350 $1.101$ 0.037-0.037 $0.059$
$\widehat{\beta}_{B}$ 0.704-0.704 $0.923$ 0.032-0.032 $0.043$ 0.009-0.009 $0.022$
Dynamic Logit Model
$\widehat{\beta}$ $0.279$ $0.332$ $0.106$ $0.119$ $0.065$ $0.073$
$\widehat{\beta}_{L}^{(1)}$ $0.101$ $0.160$ $0.026$ $0.054$ $0.011$ $0.032$
$\widehat{\beta}_{L}^{(2)}$ $0.105$ $0.164$ $0.027$ $0.055$ $0.012$ $0.032$
$\widehat{\beta}_{J}$ 2.345-2.345 $3.017$ 0.085-0.085 $0.138$ 0.020-0.020 $0.037$
$\widehat{\beta}_{B}$ 0.369-0.369 $0.407$ 0.029-0.029 $0.051$ 0.011-0.011 $0.030$
$\widehat{\rho}$ 0.546-0.546 $0.725$ 0.120-0.120 $0.169$ 0.059-0.059 $0.089$
$\widehat{\rho}_{L}^{(1)}$ 0.194-0.194 $0.396$ 0.050-0.050 $0.127$ 0.022-0.022 $0.063$
$\widehat{\rho}_{L}^{(2)}$ 0.193-0.193 $0.404$ 0.047-0.047 $0.127$ 0.019-0.019 $0.063$
$\widehat{\rho}_{J}$ 0.363-0.363 $6.295$ $0.305$ $0.577$ $0.099$ $0.267$
$\widehat{\rho}_{B}$ $0.551$ $0.734$ $0.160$ $0.186$ $0.043$ $0.071$
Dynamic Probit Model
$\widehat{\beta}$ $0.500$ $0.700$ $0.130$ $0.137$ $0.079$ $0.083$
$\widehat{\beta}_{L}^{(1)}$ $0.179$ $0.215$ $0.046$ $0.059$ $0.023$ $0.032$
$\widehat{\beta}_{L}^{(2)}$ $0.189$ $0.261$ $0.046$ $0.077$ $0.023$ $0.032$
$\widehat{\beta}_{J}$ 7.236-7.236 $9.499$ 0.499-0.499 $0.932$ 0.057-0.057 $0.104$
$\widehat{\beta}_{B}$ 2.386-2.386 $4.786$ 0.062-0.062 $0.069$ 0.018-0.018 $0.027$
$\widehat{\rho}$ 0.518-0.518 $0.781$ 0.173-0.173 $0.208$ 0.083-0.083 $0.106$
$\widehat{\rho}_{L}^{(1)}$ 0.226-0.226 $0.485$ 0.039-0.039 $0.259$ 0.024-0.024 $0.155$
$\widehat{\rho}_{L}^{(2)}$ 0.213-0.213 $0.533$ 0.037-0.037 $0.261$ 0.023-0.023 $0.167$
$\widehat{\rho}_{J}$ 4.299-4.299 $12.866$ 0.066-0.066 $2.459$ $0.116$ $1.033$
$\widehat{\rho}_{B}$ 1.621-1.621 $3.122$ $0.126$ $0.186$ $0.091$ $0.105$
@@
Table F.10: Empirical Size and Power from LR Test for Probit and Logit Models with Heterogeneous Autoregressive Coefficient, Design 1
(a) Empirical Size
Size
$(N,T)$ Uncorrected Analytical Bootstrap
Static Logit Model
$(30,30)$ $38$ $16$ 11
$(60,60)$ $41$ $11$ 33
$(90,90)$ $39$ 77 22
Static Probit Model
$(30,30)$ $78$ $41$ 99
$(60,60)$ $81$ $26$ 33
$(90,90)$ $79$ $15$ 33
Dynamic Logit Model
$(30,30)$ $62$ $37$ 88
$(60,60)$ $61$ $16$ 22
$(90,90)$ $55$ $10$ 22
Dynamic Probit Model
$(30,30)$ $93$ $78$ $52$
$(60,60)$ $93$ $45$ 77
$(90,90)$ $94$ $30$ 22
(b) Empirical Power
Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
$(N,T)$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
Static Logit Model
$(30,30)$ $83$ $44$ $13$ $37$ $36$ 88 11 11
$(60,60)$ $100$ $77$ $49$ $98$ $100$ $62$ 11 2121
$(90,90)$ $100$ $97$ $91$ $100$ $100$ $95$ 66 $91$
Static Probit Model
$(30,30)$ $99$ $85$ $12$ $27$ $61$ 1919 55 44
$(60,60)$ $100$ $99$ $53$ $99$ $100$ $87$ 0 55
$(90,90)$ $100$ $100$ $97$ $100$ $100$ $100$ 11 $88$
Dynamic Logit Model
$(30,30)$ $80$ $54$ $39$ $59$ $33$ $14$ 55 66
$(60,60)$ $100$ $77$ $59$ $99$ $98$ $42$ 11 2020
$(90,90)$ $100$ $97$ $93$ $100$ $100$ $87$ 66 $90$
Dynamic Probit Model
$(30,30)$ $99$ $91$ $67$ $71$ $81$ $62$ $44$ $40$
$(60,60)$ $100$ $98$ $76$ $100$ $100$ $81$ 22 $11$
$(90,90)$ $100$ $100$ $98$ $100$ $100$ $99$ 22 $84$
@@
F.3 Probit and Logit under Design 2
In this section, we provide the simulation results under the same settings as in Section 4, except that $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.25)$, $Z_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $Z_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$.
Table F.11: Comparisons of Uncorrected and Corrected Logit and Probit Estimates, Design 2
$(30,30)$ $(60,60)$ $(90,90)$ $(120,120)$ $(150,150)$
Bias RMSE Bias RMSE Bias RMSE Bias RMSE Bias RMSE
Static Logit Model
$\widehat{\beta}$ $0.45$ $1.20$ $0.10$ $0.13$ $0.06$ $0.07$ $0.08$ $0.05$ $0.07$ $0.04$
$\widehat{\beta}_{L}$ $0.10$ $0.20$ $0.02$ $0.06$ $0.01$ $0.04$ $0.01$ $0.03$ $0.01$ $0.02$
Static Probit Model
$\widehat{\beta}$ $0.94$ $2.28$ $0.15$ $0.25$ $0.07$ $0.08$ $0.10$ $0.05$ $0.08$ $0.04$
$\widehat{\beta}_{L}$ $0.33$ $5.17$ $0.04$ $0.06$ $0.02$ $0.04$ $0.02$ $0.03$ $0.01$ $0.02$
Dynamic Logit Model
$\widehat{\beta}$ $0.53$ $1.58$ $0.11$ $0.21$ $0.06$ $0.07$ $0.08$ $0.05$ $0.06$ $0.04$
$\widehat{\beta}_{L}^{(1)}$ $0.12$ $0.20$ $0.02$ $0.06$ $0.01$ $0.04$ $0.01$ $0.03$ $0.01$ $0.02$
$\widehat{\beta}_{L}^{(2)}$ $0.13$ $0.28$ $0.03$ $0.06$ $0.01$ $0.04$ $0.01$ $0.03$ $0.01$ $0.02$
$\widehat{\rho}$ 0.11-0.11 $0.24$ 0.06-0.06 $0.11$ 0.04-0.04 $0.07$ 0.05-0.05 $0.05$ 0.04-0.04 $0.04$
$\widehat{\rho}_{L}^{(1)}$ 0.04-0.04 $0.19$ 0.02-0.02 $0.09$ 0.01-0.01 $0.06$ 0.01-0.01 $0.04$ 0.01-0.01 $0.03$
$\widehat{\rho}_{L}^{(2)}$ 0.04-0.04 $0.20$ 0.01-0.01 $0.09$ 0.01-0.01 $0.06$ $0.00$ $0.04$ $0.00$ $0.03$
Dynamic Probit Model
$\widehat{\beta}$ $0.91$ $1.84$ $0.17$ $0.25$ $0.08$ $0.10$ $0.11$ $0.06$ $0.08$ $0.04$
$\widehat{\beta}_{L}^{(1)}$ $0.21$ $0.40$ $0.05$ $0.08$ $0.02$ $0.04$ $0.02$ $0.03$ $0.01$ $0.03$
$\widehat{\beta}_{L}^{(2)}$ $0.20$ $0.28$ $0.05$ $0.08$ $0.02$ $0.04$ $0.03$ $0.03$ $0.02$ $0.04$
$\widehat{\rho}$ 0.03-0.03 $0.16$ 0.02-0.02 $0.07$ 0.01-0.01 $0.05$ 0.02-0.02 $0.03$ 0.02-0.02 $0.03$
$\widehat{\rho}_{L}^{(1)}$ 0.02-0.02 $0.14$ 0.01-0.01 $0.07$ 0.00-0.00 $0.04$ 0.01-0.01 $0.04$ 0.01-0.01 $0.03$
$\widehat{\rho}_{L}^{(2)}$ 0.01-0.01 $0.14$ 0.00-0.00 $0.07$ 0.00-0.00 $0.05$ $0.00$ $0.04$ 0.01-0.01 $0.04$